OpenStudy (anonymous):

Prove that : \[2\sec^{2}A-\sec^{4}A-2cosec^{2}A+cosec^{4}A=\cot^{4}A-\tan^{4}A\]

OpenStudy (anonymous):

@ganeshie8 @mukushla @waterineyes

OpenStudy (anonymous):

Prove that : \[2\sec^{2}A-\sec^{4}A-2cosec^{2}A+cosec^{4}A=\cot^{4}A-\tan^{4}A\]

OpenStudy (anonymous):

\(\sec^2 A=1+ \tan^2 A \) and \( cosec^2 A=1+ \cot^2 A \) just put them in the equation and get ur answer

OpenStudy (anonymous):

i tried it but not getting

OpenStudy (anonymous):

just do it carefully

OpenStudy (klimenkov):

Use \(x^4=(x^2)^2\) and \((a+b)^2=a^2+2ab+b^2\)

OpenStudy (anonymous):

that too

OpenStudy (anonymous):

pls someone try it

OpenStudy (klimenkov):

\(\sec^4A=(\sec^2A)^2=(1+\tan^2A)^2\) Do you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i had solved it .